THE HOT BANDS OF METHANE BETWEEN 5 AND 10 MICRONS

O. Ouardi, J. C. Hilico, M. Loete, Laboratoire de Physique de l’Université de Bourgogne; and \underline{L. R. Brown}, Jet Propulsion Laboratory, California Institute of Technology, 4800 Oak Grove Drive, Pasadena, California 91109.

The experimental line intensities of 1727 transitions between 1000 and 2000 cm^{−1}belonging to the pentad-dyad system of methane are fitted to first and second order using the effective dipole moment expansion in the polyad scheme. The nine observed bands are ν_{3}-ν_{2}, ν_{3}-ν_{4}, ν_{1}-ν_{2}, ν_{1}-ν_{4}, 2ν_{4}-ν_{4}, 2ν_{2}-ν_{4}, 2ν_{2}-ν_{2}, ν_{2}+ν_{4}-ν_{2} and ν_{2}+ν_{4} -ν_{4}. The intensities are obtained from long-path spectra recorded at 0.005 cm^{−1}resolution using the McMath Fourier transform spectrometer located at Kitt Peak National Observatory. To achieve a more stable fit for three of the 27 constants required in the second order model, 1206 intensities of the dyad (ν_{2}, ν_{4}) are refitted simultaneously with the hot band data. The fits to first and second order lead to rms values, respectively, of 21.5% and 5.0% for the 1727 hot band intensities and 6.5% and 3.0% for the 1206 dyad lines. The band strengths of all 10 pentaddyad hot bands range from 0.93 for 2ν_{4}-ν_{4} to 7.7×10^{−5} for the unobserved 2ν_{4}-ν_{2} in units of cm^{−2}/atm at 296 K.

Part of the research reported in this paper was performed at the Jet Propulsion Laboratory, California Institute of Technology, under contract with the National Aeronautics and Space Administration.